OpenStudy (anonymous):

a tank contains 1200 of water. The tank is being drained at the rate of 45 gallons per minute.fine a rule for the linear function that describes the amount of water in the tank ,and use it to determine how much water will be in the tank after 15 minutes.

OpenStudy (anonymous):

your function would be Y(t)=1200-45t. the amount of water left after 15 minutes would be: Y(15)=1200-45(12)=525 gallons.